suppose that the net investment flow is described by the equestion I(t)=6t1/2 and that the initial capital stock at time t=0 is k(0) what is the time path of capital.what will be the capital formation during the time interval (3,5).

1
Expert's answer
2021-11-24T12:09:03-0500

"K(t)=\int I(t)dt=4t^\frac 32+C,"

"C=K(0),"

"K(t)=4t^\frac 32+K(0),"

"K(3,5)=4(5^\frac 32-3^\frac 32)=24."